Current spot rate of Australian dollar = $.67

Forecasted spot rate of Australian dollar 1 year from now = $.68

1-year forward rate of Australian dollar = $.93

Annual interest rate for Australian dollar deposit = 4%

Annual interest rate in the U.S. = 2%

What is your percentage return from covered interest arbitrage with $550,000 for one year?
Answer #1

For covered interest arbitrage, we need to borrow currency which has lower interest rate and invest in currency which has higher interest rate.

Step 1: borrow $550,000 for one year at 2% interest rate. amount to be returned after one-year including interest = amount borrowed*(1 + U.S. interest rate) = $550,000*(1+0.02) = $550,000*1.02 = $561,000

Step 2: convert $550,000 amount borrowed into Australian dollar using spot rate and invest at 4% for one year.

amount in Australian dollar = $550,000/$0.67 = AUD 820,895.52

Step 3: purchase a forward contract to sell Australian dollar in one year at forward rate of $0.93.

Step 4: After one year Australian dollar investment would be = AUD 820,895.52*(1+0.04) = AUD 820,895.52*1.04 = AUD 853,731.34

convert it to U.S. dollar using forward contract rate. amount in U.S. dollar = AUD 853,731.34*$0.93 = $793,970.15

profit = Investment value in U.S. dollars - amount borrowed in U.S. dollar including interest = $793,970.15 - $561,000 = $232,970.15

percentage return from covered interest arbitrage = profit/amount borrowed = $232,970.15/$550,000 = 0.4236 or 42.36%